Cove’s Cakes is a local bakery. Price and cost information follows Price per cake $ 13.31 Variable cost per cake Ingredients 2.34 Direct labor 1.05 Overhead (box, etc.) 0.22 Fixed cost per month $ 3,104.00 Required:

1. Determine Cove’s break-even point in units and sales dollars.

2. Determine the bakery’s margin of safety if it currently sells 370 cakes per month.

3. Determine the number of cubes that Cove must sell to generate $2,000 in profit.
Answer #1

1.

Break-even point (BEP): this is the point where there is no profit no loss.

BEP units = Fixed cost / Contribution margin per unit

                 = 3,104 / [Sale price – (Direct material + Direct labor + Overhead) per unit]

                 = 3,104 / [13.31 – (2.34 + 1.05 + 0.22)]

                 = 3,104 / [13.31 – 3.61]

                 = 3,104 / 9.7

                 = 320 units (Answer)

BEP sales dollar = BEP units × Sale price

                            = 320 × $13.31

                            = $4,259.20 (Answer)

2.

Margin of safety = Current sales – Breakeven sales

                                    = (370 × $13.31) – 4,259.20

                                    = 4,924.70 – 4,259.20

                                    = $665.50 (Answer)

3.

Target sales in units = (Fixed cost + Desired profit) / Contribution margin per unit

                                    = (3,104 + 2,000) / 9.70

                                    = 5,104 / 9.70

                                    = 526 units (Answer)